OpenStudy (friendzone):

What is the value of: 20 - (3 × 2^3 - 5) * 2^3 is the same as "two to the third power" or "two cubed"

OpenStudy (immanuelv):

yes it is 20 -(3 x 8 -5)

OpenStudy (immanuelv):

20-(19)

OpenStudy (anonymous):

20-3*8+5=20-24+5=1

OpenStudy (anonymous):

what we doin